# Bessel potentials and Green functions on pseudo-Euclidean spaces

### v1

@@ 3.1 General exponents–Euclidean case
Another important range of parameters is $0<{\rm Re}\mu<d$. For such $\mu$ the Bessel potential has the massless limit called the Riesz potential:
$G_{\mu,0}(x):=\lim_{m\to 0}G_{\mu,m}(x)=\frac{\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{2}\Big{)}^{\mu-d}.$ (3.8)
@@ 3.2 General exponent–antiEuclidean case
Suppose now the scalar product is negative definite. The function $\frac{1}{(-p^{2}+m^{2})^{\frac{\mu}{2}}}$ does not define uniquely a distribution, therefore one cannot compute its Fourier transform. However,
$\frac{1}{(-p^{2}+m^{2}\pm{\rm i}0)^{\frac{\mu}{2}}}:=\lim_{\epsilon\searrow 0}\frac{1}{(-p^{2}+m^{2}\pm\\ i\epsilon)^{\frac{\mu}{2}}}$ (3.9)
are tempered distributions, therefore we have two kinds of Bessel potentials:
Theorem 3.2
@@
$=\frac{{\rm e}^{\pm{\rm i}\pi\frac{d}{2}}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}U_{\frac{d-\mu}{2}}\Big{(}\frac{{\rm e}^{\pm i\pi}m^{2}|x|^{2}}{4}\Big{)}.$ (3.13)
@@
The Euclidean Bessel potential $G_{\mu,m}$ is well defined not only for $m\geq 0$, but also for ${\rm Re}(m)>0$, which guarantees $m^{2}\in{\mathbb{C}}\backslash]-\infty,0]$. Taking the limit at the imaginary line we can express the antiEuclidean Bessel potential in terms of the Euclidean one:
@@ 3.4 Averages of plane waves on sphere
Consider the sphere in ${\mathbb{R}}^{d}$ of radius $m$, denoted ${\mathbb{S}}_{m}^{d-1}={\mathbb{S}}_{m}$. Let ${\rm d}\Omega_{m}$ be the natural measure on ${\mathbb{S}}_{m}$. As an application of the Bessel potential, we will compute the Fourier transform of the measure on ${\mathbb{S}}_{m}$.
Theorem 3.3
@@ 4 General signature
Consider now a pseudo-Euclidean space of general signature ${\mathbb{R}}^{q,d-q}$. Consider two distributions regularizing $\frac{1}{(p^{2}+m^{2})^{\frac{\mu}{2}}}$.
@@
Proof of Thm 4.1. Using (A.2) and (A.4) we obtain
$\frac{1}{(2\pi)^{d}}\int\frac{{\rm e}^{{\rm i}px}{\rm d}p}{(m^{2}+p^{2}\pm{\rm i}0)^{\frac{\mu}{2}}}$
$=$ $\frac{{\rm e}^{\mp{\rm i}\frac{\pi\mu}{4}}}{(2\pi)^{d}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}t\int{\rm d}p{\rm e}^{\pm{\rm i}t(m^{2}+p^{2})}t^{\frac{\mu}{2}-1}{\rm e}^{{\rm i}px}$
$=$ $\frac{(\mp{\rm i})^{q}{\rm e}^{\mp{\rm i}\frac{\pi}{2}(\frac{\mu-d}{2})}\pi^{\frac{d}{2}}}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}t{\rm e}^{\pm{\rm i}(tm^{2}-\frac{x^{2}}{4t})}t^{\frac{\mu-d}{2}-1}.$ (4.6)
@@
For $0<{\rm Re}\mu<d$ the Bessel potentials have the massless limit called Riesz potentials:
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x):=\frac{(\pm{\rm i})^{q}\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}\pm{\rm i}0}{4}\Big{)}^{\frac{\mu-d}{2}}.$ (4.7)
@@ 5 The Minkowski signature
The Lorentzian signature is especially important, both because of its physical relevance and rich mathematical properties. The spaces ${\mathbb{R}}^{1,d-1}$ and ${\mathbb{R}}^{d-1,1}$ are two kinds of a Minkowski space, that is, a pseudo-Euclidean space with a Lorentzian signature. We will treat ${\mathbb{R}}^{1,d-1}$ as the standard form of the Minkowski space. $x^{0}$ will denote the first coordinate of ${\mathbb{R}}^{1,d-1}$, which we assume to be timelike (having a negative coefficient in the scalar product). The remaining, spacelike coordinates will be denoted $\vec{x}$, so that $x=(x^{0},\vec{x})$. In other words,
@@ 5.1 General exponent
Proof.  For definiteness, consider (5.4) with the minus sign. In order to prove that it is contained in $J^{\vee}$, by the Lorentz invariance it suffices to prove that it is zero for $x^{0}<0$. We write
@@
(Let us explain the correctness of the formula (5.7) for the advanced and retarded Bessel potential. We have $d-\mu<d$, and therefore we can apply Proposition 1 below, which shows that we can multiply the distribution $\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}$ by the discontinuous function $\theta(\pm x^{0})$. The resulting distribution is then multiplied by the smooth function ${\bf F}_{\frac{\mu-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}$.)
@@
where again we used the notation introduced in (2.47). (5.9) is clearly supported in $J^{\wedge}\cup J^{\vee}$. By Thm 5.1, we know that $G_{\mu,m}^{\vee/\wedge}$ are supported in $J^{\vee/\wedge}$. Thus to find expressions for $G_{\mu,m}^{\vee/\wedge}$ we need to “split the distribution” (5.9) into two terms, one supported in $J^{\vee}$ and the other in $J^{\wedge}$.
Using Proposition 1. to justify the multiplication of a distribution (5.9) by the (discontinuous) function $\theta(\pm{x}^{0})$, we can define
@@
so $c_{\alpha,m}^{\vee/\wedge}=0$ for even $|\alpha|$. $G_{\mu,m}^{\vee/\wedge}$ and $\tilde{G}_{\mu,m}^{\vee/\wedge}$ are invariant with respect to the action of the connected component of the Lorentz group. The same must apply to their difference $B_{\mu,m}^{\vee/\wedge}$. Derivatives evaluated at 00 transform as vectors under the action of the Lorentz group. However, $\langle B_{\mu,m}^{\vee/\wedge},\phi\rangle$ is a sum of terms with only odd number of indices so it cannot be invariant under the action of the Lorentz group unless $B_{\mu,m}^{\vee/\wedge}=0$. We conclude that $G_{\mu,m}^{\vee/\wedge}=\tilde{G}_{\mu,m}^{\vee/\wedge}$. $\Box$
@@
Below we will prove a technical result implying that the expression (5.10) yields a well-defined distribution. We will use the identification ${\mathbb{R}}^{1,d-1}\simeq{\mathbb{R}}\times{\mathbb{R}}^{d-1}$, the first variable denoted $x^{0}$ or $t$. For $n,m\in{\mathbb{N}}_{0}$ and $\chi\in{\mathcal{S}}({\mathbb{R}}^{1,d-1})$ let us introduce the semi-norms, which involve only the variables $\vec{x}\in{\mathbb{R}}^{d-1}$:
@@
$\Big{|}\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}\phi(x){\rm d}x\Big{|}\leq\sum_{k=0}^{\lfloor\frac{{\rm Re}\nu}{2}\rfloor}\int c_{k}|t|^{d-{\rm Re}\nu+k-1}\|\phi(t,\cdot)\|_{0,k}{\rm d}t,$ (5.14)
@@
For $0<{\rm Re}\mu<d$ we also have the massless Riesz potentials:

### v2

@@ 3.1 General exponents–Euclidean case
3.2 General exponents–massless case
For $0<{\rm Re}\mu<d$ the following function is in $L_{\rm loc}^{1}({\mathbb{R}}^{d})$ and is bounded at infinity, hence it defines a regular distribution in ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$:
$G_{\mu,0}(x)$ $:=\int\frac{{\rm e}^{{\rm i}px}}{|p|^{\mu}}\frac{{\rm d}p}{(2\pi)^{d}}$ (3.8)
$=\frac{\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{|x|}{2}\Big{)}^{\mu-d}.$ (3.9)
It is called the Riesz potential, and it is the massless limit of Bessel potentials:
Theorem 2
Let $0<{\rm Re}\mu<d$. Then
$G_{\mu,0}(x)=\lim_{m\to 0}G_{\mu,m}(x)$ (3.10)
in the sense of ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$.
Proof.  One can prove this fact in the position space, see Subsection 4.2, where we give a proof in the case of a general signature. Instead, in this section we describe a proof based on the momentum space.
For $0<\mu<d$, $|p|^{-\mu}$ is a regular distribution. By using the Dominated Convergence Theorem we see that the pointwise limit
$\lim_{m\to 0}(p^{2}+m^{2})^{-\frac{\mu}{2}}=|p|^{-\mu}$ (3.11)
is a limit in the sense of ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$. The Fourier transformation is a continuous operator on ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$. Therefore, for considered $\mu$, (3.10) is true. $\Box$
@@ 3.3 General exponent–antiEuclidean case
Suppose now the scalar product is negative definite. For $m^{2}>0$, the function $\frac{1}{(-p^{2}+m^{2})^{\frac{\mu}{2}}}$ does not define uniquely a distribution, therefore one cannot compute its Fourier transform. However, if $m^{2}\in{\mathbb{C}}\backslash[0,\infty[$, then $\frac{1}{(-p^{2}+m^{2})^{\frac{\mu}{2}}}$ is a tempered distribution, and one can take its limit from above or below in the distributional sense:
$\frac{1}{(-p^{2}+m^{2}\pm{\rm i}0)^{\frac{\mu}{2}}}:=\lim_{\epsilon\searrow 0}\frac{1}{(-p^{2}+m^{2}\pm{\rm i}\epsilon)^{\frac{\mu}{2}}}.$ (3.12)
Thus we obtain two kinds of Bessel potentials in the antiEuclidean case:
Theorem 3
@@
$=\frac{{\rm e}^{\pm{\rm i}\pi\frac{d}{2}}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}U_{\frac{d-\mu}{2}}\Big{(}\frac{{\rm e}^{\pm{\rm i}\pi}m^{2}x^{2}}{4}\Big{)}.$ (3.16)
@@
Note that the Euclidean Bessel potential $G_{\mu,m}$ is well defined not only for $m\geq 0$, but also for ${\rm Re}(m)>0$, which guarantees $m^{2}\in{\mathbb{C}}\backslash]-\infty,0]$. Taking the limit at the imaginary line we can express the antiEuclidean Bessel potential in terms of the Euclidean one:
@@ 3.5 Averages of plane waves on sphere
Consider the sphere in ${\mathbb{R}}^{d}$ of radius $m$, denoted ${\mathbb{S}}_{m}^{d-1}={\mathbb{S}}_{m}$. Let ${\rm d}\Omega_{m}$ be the natural measure on ${\mathbb{S}}_{m}$. As an application of Bessel potentials, we will compute the Fourier transform of the measure on ${\mathbb{S}}_{m}$.
Theorem 4
@@ 4 General signature
4.1 Positive mass
Consider now a pseudo-Euclidean space of general signature ${\mathbb{R}}^{q,d-q}$. $\frac{1}{(p^{2}+m^{2})^{\frac{\mu}{2}}}$ no longer defines a tempered distribution in the general signature. Just as in the antiEuclidean case, there are two natural regularizations of this function:
@@ 4.1 Positive mass
Proof of Thm 5. Using (A.2) and (A.4) we obtain
$\frac{1}{(2\pi)^{d}}\int\frac{{\rm e}^{{\rm i}px}{\rm d}p}{(m^{2}+p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}$
$=$ $\frac{{\rm e}^{\pm{\rm i}\frac{\pi\mu}{4}}}{(2\pi)^{d}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}t\int{\rm d}p{\rm e}^{\mp{\rm i}t(m^{2}+p^{2})}t^{\frac{\mu}{2}-1}{\rm e}^{{\rm i}px}$
$=$ (±$\frac{(\pm{\rm i})^{q}{\rm e}^{\pm{\rm i}\frac{\pi}{2}(\frac{\mu-d}{2})}\pi^{\frac{d}{2}}}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{\mu}{2})}\int_{0}^{\infty}{\rm d}t{\rm e}^{\mp{\rm i}(tm^{2}-\frac{x^{2}}{4t})}t^{\frac{\mu-d}{2}-1}.$ (4.6)
@@
4.2 Zero mass
For $0<{\rm Re}\mu<d$ let us introduce two distributions in ${\mathcal{S}}^{\prime}$
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ $:=\int\frac{{\rm e}^{{\rm i}px}}{(p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}\frac{{\rm d}p}{(2\pi)^{d}}$ (4.7)
=(±$=\frac{(\pm{\rm i})^{q}\Gamma(\frac{d-\mu}{2})}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Big{(}\frac{x^{2}\pm{\rm i}0}{4}\Big{)}^{\frac{\mu-d}{2}}.$ (4.8)
They will be called Feynman/antiFeynman Riesz potentials. They are massless limits of the corresponding Bessel potentials:
Theorem 7
For $0<{\rm Re}\mu<d$ we have
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x)=\lim_{m\searrow 0}G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}(x)$ (4.9)
in the sense of ${\mathcal{S}}^{\prime}$.
Proof.  Surprisingly, a momentum space proof, from the Euclidean case, seems to be difficult to generalize to the non-Euclidean case. Instead, we will present a proof in the position space.
Using the decomposition (2.45) of the function $U_{\alpha}$, we can write
$G_{\mu,m}(x)$ =(±$=\frac{(\pm{\rm i})^{q}}{\Gamma(\frac{\mu}{2})(4\pi)^{\frac{d}{2}}}\Bigg{(}\sum_{j=0}^{\lfloor{\rm Re}\frac{d-\mu}{2}\rfloor-1}\frac{(-1)^{j}m^{j}\Gamma(\frac{d-\mu}{2}-j)}{j!}\Big{(}\frac{(x^{2}\pm{\rm i}0)}{4}\Big{)}^{j-\frac{d-\mu}{2}}$ (4.10)
$+m^{d-\mu}U_{\frac{d-\mu}{2}}^{\rm reg}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}\Bigg{)}.$ (4.11)
The line (4.10) obviously converges to (4.8). By (2.7), $U_{\frac{d-\mu}{2}}^{\rm reg}$ is a continuous function of a polynomial growth at infinity. Therefore, the second line (4.11) converges to zero in ${\mathcal{S}}^{\prime}$. $\Box$
Note that as a consequence of the above theorem and of the continuity of the Fourier transformation on ${\mathcal{S}}^{\prime}({\mathbb{R}}^{d})$ we can infer that
$\lim_{m\searrow 0}\frac{1}{(p^{2}+m^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}=\frac{1}{(p^{2}\mp{\rm i}0)^{\frac{\mu}{2}}}$ (4.12)
in the sense of ${\mathcal{S}}^{\prime}$.
4.3 Scaling degree of distributions
Let us start by defining the action of a dilation by $\lambda$ on a distribution $T(x)$ as $T_{\lambda}(x)=T(\lambda x)$, by which we mean the dual action to the dilation on test functions
$\langle T_{\lambda}|f\rangle=\int T(\lambda x)f(x){\rm d}x=\lambda^{-d}\int T(x)f(\lambda^{-1}x){\rm d}x.$ (4.13)
Given a distribution $T\in\mathcal{D}^{\prime}({\mathbb{R}}^{d})$, we define its scaling degree $\mathrm{sd}(T)$ as
$\mathrm{sd}(T)=\inf\Big{\{}\omega:\lim_{\lambda\searrow 0}\lambda^{\omega}T_{\lambda}=0\ \text{ in }\mathcal{D}^{\prime}({\mathbb{R}}^{d})\Big{\}}.$ (4.14)
The scaling degree of a distribution is often used in mathematical analysis of Quantum Field Theory [6, 7].
Let us compute the scaling degree of Bessel potentials.
Theorem 8
$\mathrm{sd}G_{m,\mu}^{\mathrm{F}/\overline{\mathrm{F}}}=\begin{cases}d-\mu,&0<\mu\leq d;\\ 0,&d\leq\mu.\end{cases}.$ (4.15)
Proof.  For $0<\mu<d$, the Riesz potentials $G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}$ defined in (4.8) are homogeneous:
$G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(\lambda x)=\lambda^{\mu-d}G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}(x).$ (4.16)
So $\mathrm{sd}G_{\mu,0}^{\mathrm{F}/\overline{\mathrm{F}}}=d-\mu$.
By the definition of the Bessel potential, the mass dependence is (1.2)
$G_{\mu,m}(\lambda x)=\lambda^{\mu-d}G_{\mu,\lambda m}(x),$ (4.17)
so, according to Theorem 7,
$\lim_{\lambda\searrow 0}\lambda^{d-\mu}G_{\mu,m}(\lambda x)=\lim_{\lambda\searrow 0}G_{\mu,\lambda m}(x)=G_{\mu,0}(x),$ (4.18)
which shows that $\mathrm{sd}G_{\mu,m}^{\mathrm{F}/\overline{\mathrm{F}}}=d-\mu$ for any mass $m$ and $0<\mu<d$.
For $d<\mu$, $G_{m}^{F/\overline{\mathrm{F}}}$ is a continuous bounded function, so its scaling degree is 00.
For $d=\mu$, we have
$G_{d,m}(x)=\frac{(\pm{\rm i})^{q}\sqrt{\pi}m^{d-\mu}}{\Gamma(\frac{d}{2})(4\pi)^{\frac{d}{2}}}U_{0}\Big{(}\frac{m^{2}(x^{2}\pm{\rm i}0)}{4}\Big{)}.$ (4.19)
Now, we can use the bound (2.7) and the expansion (2.1)
[…]
where $H$ is an entire function, just as $\mathbf{F}_{0}$. Using this we easily show that for $\omega>0$
$\lambda^{\omega}G_{d,m}(\lambda x)\to 0$ (4.22)
in the sense of ${\mathcal{S}}^{\prime}$. $\Box$
@@ 5 The Minkowski signature
The Lorentzian signature is especially important, both because of its physical relevance and rich mathematical properties. The spaces ${\mathbb{R}}^{1,d-1}$ and ${\mathbb{R}}^{d-1,1}$ are two kinds of a Minkowski space, that is, a pseudo-Euclidean space with a Lorentzian signature. We will treat ${\mathbb{R}}^{1,d-1}$ as the standard form of a Minkowski space. $x^{0}$ will denote the first coordinate of ${\mathbb{R}}^{1,d-1}$, which we assume to be timelike (having a negative coefficient in the scalar product). The remaining, spacelike coordinates will be denoted $\vec{x}$, so that $x=(x^{0},\vec{x})$. In other words,
@@ 5.1 General exponent
Proof.  For definiteness, consider (5.4) with the minus sign. In order to prove that its support is contained in $J^{\vee}$, by the Lorentz invariance it suffices to prove that it is zero for $x^{0}<0$. We write
@@
Formula (5.7) involves the multiplication of a distribution by a discontinuous function, which in general is not well defined. At the end of this subsection we provide explain how this formula can be correctly interpreted.
@@
where again we used the notation introduced in (2.47). (5.9) is clearly supported in $J^{\wedge}\cup J^{\vee}$. By Thm 9, we know that $G_{\mu,m}^{\vee/\wedge}$ are supported in $J^{\vee/\wedge}$. Thus to find expressions for $G_{\mu,m}^{\vee/\wedge}$ we need to “split the distribution” (5.9) into two terms, one supported in $J^{\vee}$ and the other in $J^{\wedge}$.
Using Proposition 11 to justify the multiplication of a distribution (5.9) by the (discontinuous) function $\theta(\pm{x}^{0})$, we can define
@@
so $c_{\alpha,m}^{\vee/\wedge}=0$ for even $|\alpha|$. $G_{\mu,m}^{\vee/\wedge}$ and $\tilde{G}_{\mu,m}^{\vee/\wedge}$ are invariant with respect to the action of the proper Lorentz group. The same must apply to their difference $B_{\mu,m}^{\vee/\wedge}$. Derivatives evaluated at 00 transform as vectors under the action of the Lorentz group. However, $\langle B_{\mu,m}^{\vee/\wedge},\phi\rangle$ is a sum of terms with only odd number of indices so it cannot be invariant under the action of the Lorentz group unless $B_{\mu,m}^{\vee/\wedge}=0$. We conclude that $G_{\mu,m}^{\vee/\wedge}=\tilde{G}_{\mu,m}^{\vee/\wedge}$. $\Box$
@@
For $0<{\rm Re}\mu<d$ we also have the massless Riesz potentials:
As we mentioned above, the formula (5.7) for the advanced and retarded Bessel potential involves a product of two distributions, and therefore it needs a justfication. We will explain two approaches how to interpret this formula.
The first approach is quite elementary. It uses the identification ${\mathbb{R}}^{1,d-1}\simeq{\mathbb{R}}\times{\mathbb{R}}^{d-1}$, with the first variable denoted $x^{0}$ or $t$. For the remaining variables $\vec{x}$ we will later use spherical coordinates $(r,\Omega)$ with $r=|\vec{x}|$. For $n,m\in{\mathbb{N}}_{0}$ and $\chi\in{\mathcal{S}}({\mathbb{R}}^{1,d-1})$ let us introduce the semi-norms, which involve only the variables $\vec{x}\in{\mathbb{R}}^{d-1}$:
@@
$\Big{|}\int\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}\phi(x){\rm d}x\Big{|}\leq\sum_{k=0}^{\lfloor\frac{{\rm Re}\nu}{2}\rfloor}\int c_{k}|t|^{d-{\rm Re}\nu+k-1}\|\phi(t,\cdot)\|_{0,k}{\rm d}t,$ (5.16)
@@
Now we have $d-\mu<d$, and therefore Proposition 11 shows that we can multiply the distribution $\big{(}\tfrac{x^{2}}{4}\big{)}_{-}^{\frac{\mu-d}{2}}$ by the discontinuous but bounded function $\theta(\pm x^{0})$. The resulting distribution is then multiplied by the smooth function ${\bf F}_{\frac{\mu-d}{2}}\Big{(}\frac{m^{2}x^{2}}{4}\Big{)}$, obtaining the right hand side of (5.7).
An alternative way to define the product in (5.7) is based on the concept of the wave front set [20]. Here are the wave front sets of the distributions contained in (5.7):
$\mathrm{WF}\left(\theta(t)\right)=$$\Big{\{}\left((0,\vec{x}),(\tau,0)\right)\ :\ \vec{x}\in\mathbb{R}^{d-1},\tau\neq 0\Big{\}},$
$\mathrm{WF}\left(\big{(}x^{2}\big{)}_{-}^{-\frac{\nu}{2}}\right)=$$\Big{\{}\left((t,\vec{x}),(-\lambda t,\lambda\vec{x})\right)\ :\ t^{2}-\vec{x}^{2}=0,(t,\vec{x})\neq 0,\lambda\neq 0\Big{\}}$
$\cup\Big{\{}((0,0),(\tau,\vec{k}))\ :\ \tau^{2}-\vec{k}^{2}=0,(\tau,\vec{k})\neq 0\Big{\}},$
where $(\tau,\vec{k})$ denotes the dual variable to $(t,\vec{x})$. The fiberwise sum of wavefront sets $\mathrm{WF}\left(\theta(t)\right)+\mathrm{WF}\left(\left(x^{2}\right)_{-}^{-\frac{\nu}{2}}\right)$ does not contain an element of the form $((t,\vec{x}),(0,0))$. Therefore, by Hörmander’s criterion [20, p. 267], the product of these two distributions is well defined.